A flux quantum (fluxoid) is approximately equal to 2×10^-7 gauss−cm2.A type II superconductor is placed in a small magnetic field, which is then slowly increased till the field starts penetrating the superconductor. The strength of the field at this point is 2/π×10^5 gauss

The penetration depth of this superconductor is

 100 Å

 200 Å

 1000 Å

 314 Å


1
Expert's answer
2021-10-15T10:23:47-0400

The relation between critical field and penetration depth is given by


"H_c=\frac{\Phi_0}{\pi\lambda^2}"

Hence, the penetration depth

"\lambda=\sqrt{\frac{\Phi_0}{\pi H_c}}"

"=\sqrt{\frac{2×10^{-7}\:\rm gauss*cm^2}{\pi *2/π×10^5\:\rm gauss}}=10^{-6}\:\rm cm=100\:\mathring{A}"

Answer: "100\rm\:\mathring{A}"